### v2

@@ IV.4 Assessing student misconceptions
Wilson et al. [29] and Harrison et al. [30] describe the concept of uncertainty of physical measurements as a threshold concept in physics, emphasizing its critical role in deepening students’ understanding of various topics, including experimental data analysis and the fundamental principles of quantum physics [31]. Majidi’s review article further highlights that misconceptions about uncertainties can arise from three primary sources: the knowledge itself, teaching methods, or the students’ perspectives. Overall, misconceptions can be more detrimental than gaps in knowledge. This is because students tend to resist new information that conflicts with their existing beliefs, resulting in cognitive dissonance and frustration. Our ability to characterize misconceptions through the CDPA however is constrained to assessing the alternatives present in the test and mapping it to possible misconceptions.
@@
Furthermore, Wilson et.al [29] discusses that understanding uncertainties can be divided into three different layers of understading: pattern recognition, formal procedure, and meaning. In this study, three skills were evaluated through interviews with students, and it was discussed that integrating the three elements can give a more complete view on student understanding. In terms of the CDPA, pattern recognition is applicable to questions that evaluate uncertainties in graphs, or numerical uncertainties. The second skill, formal procedure, is tested when students have to actually calculate the uncertainties for some of the numerical questions presented in the assessment. The skill related to the meaning of uncertainties cannot be directly assessed through the CDPA, but can potentially be inferred through the experimental outcomes of the laboratory activities.
@@
By analyzing the results from novice to expert levels in two Canadian universities (Figure 1), we see that the CDPA is not an easy test. The distractors represented by each of the incorrect answers, which are linked to distinct numerical values and graphical representations, were not selected at random. Rather, they reflect a specific approach to reasoning that students could easily be led astray when answering the ten questions. Thus, in light of the characteristic challenges of this scheme, instead of interpreting the mean CDPA score as a final measurement, we propose a different approach to these results by focusing on the distribution of correct and incorrect responses for each question. We use this distribution as an additional layer of analysis that can be derived when using this instrument. The original CDPA paper divides the survey by specific abilities and classifies them in two distinct groupings. The first grouping, composed of Q1, Q2, Q5, Q6, Q9, and Q10 evaluates students’ understanding of the meaning and use of uncertainty in measurements. The second grouping, Q3, Q4, Q7 and Q8, measures their ability to relate functions, graphs, and numbers. In the analysis that follows, we chose to focus our attention on the first grouping in the analysis that follows since these questions specifically target the practical understanding of uncertainties. For further reference, a list of the ten questions, complete with their distractor answers and the justification for these answers, is available in the original paper [10].
@@
In Figure 4, we present our results taken from PHYS 257 and PHYS 258 for two complete years of undergraduate experimental physics laboratory education. The pie charts show the correct answers highlighted in green and the incorrect answers represented in gray. Remarkably, we notice that the two different cohorts displayed in Figure 4, over two distinct years, show a very similar distribution of responses. This trend becomes particularly evident when narrowing in on the correct responses given by respondents from both PHYS 257 and PHYS 258. When comparing the wrong alternative answers one by one, we also identify similarities that might point to distinct approaches to student reasoning. Moreover, in comparing the progression from one course to another, we notice a positive increase towards the correct answer for most questions.
@@
The professors we reported our findings to, each of whom had instructed the courses included in this study, were asked to analyze the descriptions and justifications of the detractor questions and concluded that they were in agreement with the parameters set by the original CDPA paper. After being presented with the range of answers gathered from their respective students, they were then prompted to evaluate the content of each incorrect answer to further examine the misconceptions associated with each possible alternative approach to reasoning. The instructors were then asked to classify whether students should be able to answer each question by the end of the course they instructed. Their responses were categorized as follows: 1) yes, these students should be able to answer the question, 2) no, these students should not be able to answer the question, or 3) maybe, a stronger student in this course might be able to answer the question but it is not a requirement to pass the course in question. The analysis for each question, indicated in Table 3 below, reflects the results of these feedback discussions. Overall, the mapping of misconceptions established by these McGill professors is in line with the previously published results.
@@
Figure 5: CDPA fraction of correct responses for each of the 10 questions for PHYS 359 in 2022.
We make a question by question analysis of the results, presented in Figure 5. When considering Q1, for example, students are still being exposed to the content and do not necessarily need to know how to answer the CDPA questions after completing PHYS 257. The only exception to this was possibly for a student with a stronger background of knowledge. The professors also explained that the incorrect alternatives show that students disregard uncertainties when taking an average value. There was further consensus that, by the end of PHYS 258, students should be able to identify and treat different uncertainties in measurements correctly.
Q2 presents the highest correct response rate of all the CDPA questions for this student demographic before the start of PHYS 257. This question addresses ability to understand significant figures in astronomical units. According to the professors, students should have entered the course already able to respond to this question. Incorrect alternatives do not present any specific insight into student reasoning other than a misconception on the expression of significant figures.
In contrast, the professors pointed out that only students who are very strong should be able to answer Q5 and Q6, and that their understanding of the concepts in these questions are presented and explored throughout PHYS 258. However, students are still not ready by the end of the course to compare different models for best fit, an approach they are only exposed to more in depth at a higher level. It must be reiterated that all the CDPA distractor answers are very close to the correct response and that, for these two questions, choosing an incorrect answer can be associated with disregarding numerical uncertainties graphically when considering a best fit line. For example, students often choose to fit the most points rather than fit the ones with the smallest uncertainty, an approach to reasoning that is indicative of their lack of knowledge on how to prioritize.
Most students answer Q9 correctly and they should be able to respond to it only by the end of PHYS 258 because, during PHYS 257, they are still being exposed to the content. In this case, the alternatives that are incorrect reflect a misunderstanding of statistics and how statistics impact experimental measurements.
Q10 investigates students’ capacity to combine numerical uncertainties from different measurements. The professors explain that they expect students to master this critical ability by the end of PHYS 258, as the students develop a better understanding of how to handle uncertainties through the progression of experiments they conduct during the course, culminating in a final project they complete.
@@ IV.5 What the CDPA reveals about upper level lab courses
Having established a clearer picture of what the CDPA reveals about students’ understanding and misunderstanding of concepts related to uncertainty during their first comprehensive year of introductory lab courses, we now turn to what the CDPA can tell us about the level of learning that takes place in more advanced laboratory courses. As previously explained, beyond the 200-level courses, this study also includes data collected at the 300- and 400-levels, most notably PHYS 359 - Advanced Physics Laboratory 1, reserved for Honors Physics undergraduate students, and PHYS 439 - Majors Laboratory in Modern Physics, the equivalent course for Physics majors.
@@
Between 2019 and 2022, the high variance in CDPA scores for PHYS 258 demonstrates that, at this point in their learning, students still do not have a complete mastery of the skills needed to evaluate experimental uncertainties. For PHYS 359, however, we note a significant increase in the CDPA score in 2022 that warrants a more in-depth investigation into what factors may have contributed to this leap in student understanding. As can be seen in Figure 2, the 2022 mean CDPA score was significantly higher, from a class average of $2.5\pm 2.0$ by the end of PHYS 258 to $4.5\pm 1.0$ by the end of PHYS 359. Although we can expect noticeable improvement in the CDPA score after this upper-level laboratory course, the extent of this statistically significant increase over one semester left us with more questions.
@@
Through discussion with the instructor who taught PHYS 359 during the Winter 2022 semester, we could ascertain that certain structural elements of the course contributed to this demonstrable increase in student understanding of uncertainties. Three key factors stand out in the methods of the instruction and assessment of that semester. In terms of how the course was taught, the most significant difference between PHYS 258 and PHYS 359 is the class size. While PHYS 258 had over 100 students, the smaller enrollment numbers in PHYS 359 allowed for greater one on one interaction between students and their instructor. Furthermore, students also spend more time in the lab per week, and perform experiments that require a higher degree of precision and attention to uncertainties.
@@
The course has been going through changes and adaptations. One example is that a course instructor for PHYS 359 in 2022 identified a noticeable weakness in incoming students’ level of understanding of uncertainties. This led to the development and implementation of an additional module dedicated to curve fitting in a self-paced Python tutorial. This module offered students more advanced exercises emphasizing graph interpretation and analysis, incorporating an interactive simulation feature and enabling personalized feedback. The teaching tool allowed the instructor to leverage the smaller class size for a more individualized instructional approach, dedicating time to helping students apply concepts related to uncertainties. With this personalized tool, the instructor could identify and correct student misconceptions more precisely.
@@ V Conclusions
By comparing average CDPA scores from pre and post tests across all core laboratory courses in the McGill experimental physics undergraduate curriculum, we were able to distinguish distinct patterns in students’ understanding of concepts related to uncertainties from introductory 100-level courses through to upper-level 300-level and 400-level courses. As a general observation, our results show that, at the 100-level, students’ CDPA scores remained quite low and were mostly on-par with random guesswork. These results confirm UBC’s findings for introductory-level courses, which can be explained by the fact that, in first-year courses, students are introduced to the concept of uncertainties for the first time and are only expected to solve simple mechanics problems and conduct experiments without any deeper conceptual understanding.
@@
At the 200-level, however, we notice an increase in students’ CDPA scores that remained largely consistent across all four years of this study. This uptick in student understanding can be explained by the fact that it is during this full academic year of laboratory courses that students learn how to perform curve fitting, start considering uncertainties to assess the outcome of an experiment, and produce and interpret graphs based on the measurement of uncertainties. Whereas in PHYS 257, students are asked to assess systematic errors and perform error propagation, in PHYS 258, they need to consider statistical uncertainties to get the best-fit curves, understand and justify their reasoning, and differentiate between different models. There is additional pedagogical value in the fact that PHYS 258 culminates in a small research project in which students are required to answer a research question of their choice, experimentally validate their measurements and demonstrate understanding of how the uncertainties should be considered in the analysis. Nevertheless, it is not until the 300-level and 400-level courses, when 20 percent of the lab report grade is allocated to justifying quantitative statistical uncertainties, that students’ demonstrated understanding of uncertainties develops into holistic concepts. In addition to using the CDPA instrument as a means to rank student understanding across the departmental laboratory curricula, our research also shows that this tool can be used to help instructors comprehend the nature of their students’ misconceptions. As shown in the breakdown of CDPA incorrect responses, Figure 4, CDPA results can be used as a remediation tool to help professors identify gaps in student understanding, thereby providing an additional means to refine their teaching approach for subsequent cohorts.
@@
In conclusion, through this collaborative study, we were able to highlight new opportunities for pedagogical and structural improvement. We show how a better understanding of students misconceptions can be used enhances the professors’ understanding on student learning and improvements in a course-by-course basis. This research reinforces the value of collecting program level data and reporting the findings back to course professors for discussion. Furthermore, without engaging the professors in the process of data analysis, we would not have been able to pinpoint an educational benefit featured in this research, which can be tied to a simulation exercise developed by one professor that allows students to individually probe graphing with uncertainties. This type of resource, which could be adapted to pinpoint the learning needs of students across different levels of laboratory courses, shows promise to impact students’ conceptual understanding of uncertainties in addition to the existing course curriculum. Even without any additional resources, this study also points to a potential need for more inquiry-guided learning in laboratory courses, or, at the very least, it serves as an impetus to carve out more personalized attention in large-scale introductory laboratory courses.

### v3

@@ IV.4 Assessing student misconceptions
Wilson et al. [29] and Harrison et al. [30] describe the concept of uncertainty of physical measurements as a threshold concept in physics, emphasizing its critical role in deepening students’ understanding of various topics, including experimental data analysis and the fundamental principles of quantum physics [31]. Majidi’s review article further highlights that misconceptions about uncertainties can arise from three primary sources: the knowledge itself, teaching methods, or the students’ perspectives. Overall, misconceptions can be more detrimental than gaps in knowledge. This is because students tend to resist new information that conflicts with their existing beliefs, resulting in cognitive dissonance and frustration. We can use the CDPA to identify misconceptions, acknowledging that we are limited to assessing the alternatives present in the test.
@@
Furthermore, Wilson et.al [29] discuss that understanding uncertainties can be divided into three different layers of understanding: pattern recognition, formal procedure, and meaning. In this study, three skills were evaluated through interviews with students, and it was discussed that integrating the three elements can give a more complete view on student understanding. In terms of the CDPA, pattern recognition is applicable to questions that evaluate uncertainties in graphs, or numerical uncertainties. The second skill, formal procedure, is tested when students have to actually calculate the uncertainties for some of the numerical questions presented in the assessment. The skill related to the meaning of uncertainties cannot be directly assessed through the CDPA, but can potentially be inferred through the experimental outcomes of the laboratory activities.
@@
By analyzing the results from novice to expert levels in two Canadian universities (Figure 1), we see that the CDPA is not an easy test. The distractors represented by each of the incorrect answers, which are linked to distinct numerical values and graphical representations, were not selected at random. The distractors reflect a faulty approach to reasoning that students might use and that might lead them astray. Instead of interpreting the mean CDPA score as a final measurement, we propose a different approach wherein we focus on the distribution of correct and incorrect responses. We use this distribution as an additional layer of analysis that can be derived when using this instrument. The original CDPA paper divides the survey by specific abilities and classifies them in two distinct groupings. The first grouping, composed of Q1, Q2, Q5, Q6, Q9, and Q10 evaluates students’ understanding of the meaning and use of uncertainty in measurements. The second grouping, Q3, Q4, Q7 and Q8, measures their ability to relate functions, graphs, and numbers. In the analysis that follows, we chose to focus our attention on the first grouping since these questions specifically target the practical understanding of uncertainties. For further reference, a list of the ten questions, complete with their distractor answers and the justification for these answers, is available in the original paper [10].
@@
In Figure 4, we present the results for two complete years of undergraduate experimental physics laboratory education, PHYS 257 and PHYS 258. The pie charts show the correct answers highlighted in green and the incorrect answers are represented in gray. Remarkably, we notice that the two cohorts displayed in Figure 4 show a very similar distribution of responses despite having taken the test in two different years. This trend becomes particularly evident when narrowing in on the correct responses given by respondents from both PHYS 257 and PHYS 258. When comparing the wrong alternative answers one by one, we also identify similarities that might point to distinct approaches to student reasoning. Moreover, in comparing the progression from one course to another, we notice a positive increase towards the correct answer for most questions.
@@
We met with the professors to review the CDPA test questions. The professors agreed that the distractor responses indicated similar misconceptions to those described in the original CDPA paper. The professors evaluated each incorrect answer to identify possible misconceptions underlying the error. The professors were then asked to classify whether students should be able to answer each question by the end of the course they instructed. Their responses were categorized as follows: 1) yes, students should be able to answer the question, 2) no, students should not be able to answer the question, or 3) maybe, a stronger student in this course might be able to answer the question but it is not a requirement to pass the course in question. The analysis for each question, indicated in Table 3 below, reflects the results of these feedback discussions. Overall, the mapping of misconceptions established by these McGill professors is in line with the original CDPA paper.
@@
Figure 5: Distribution of correct responses for all the 10 CDPA questions for students in PHYS 131, PHYS 257, PHYS 258, and PHYS 359. The stars represent the degree of statistical significance in these results.
We present the results of individual questions in Figure 5. We notice an overall trend of increasing correct answers as students progress through their degree. To better understand these results, in Table 3, we show the possible misconceptions influencing the students’ responses in the core experimental physics courses, focusing on PHYS 257 and PHYS 258. Students in their early years respond to Q1 incorrectly which is expected since they do not yet have the necessary skillset. The professors noted that the incorrect alternatives in Q1 show that students disregard uncertainties when taking the average of experimental data points. However they agreed that by the end of PHYS 258, students should be able to identify and treat different uncertainties in measurements correctly.
Q2 received the highest rate of correct responses by students in all courses. This question addresses ability to understand significant figures in astronomical units. According to the professors, students should be able to respond to this question even before beginning the course. Incorrect alternatives do not present any specific insight into student reasoning other than a possible misconception about the expression of significant figures.
In contrast, the professors pointed out that only students in PHYS 257 and PHYS 258 who are very strong should be able to answer Q5 and Q6. The concepts in these questions are introduced in PHYS 258, however students are not expected to compare different models for best fit until upper level courses. It must be reiterated that all the CDPA distractors are very close to the correct response and that, for these two questions, choosing an incorrect answer can be associated with disregarding numerical uncertainties graphically when considering a best fit line. For example, students often choose to fit the most points rather than fit the ones with the smallest uncertainty, an approach to reasoning that is indicative of their lack of knowledge on how to prioritize.
Q9 evaluates how students deal with the outcome of measurements and associated instrumental uncertainties. Students are not expected to answer this question correctly until they reach the end of PHYS 258. The distractors may reflect a misunderstanding of statistics and how statistics impact experimental measurements.
Q10 investigates students’ capacity to combine numerical uncertainties from different measurements. The professors expect students to master this critical ability by the end of PHYS 258, as the students develop a better understanding of how to handle uncertainties through the progression of experiments they conduct during the course, culminating in a final project they complete. This is reflected in the results when comparing PHYS 258 with earlier years.
@@ IV.5 What the CDPA reveals about upper level lab courses
Having established a clearer picture of what the CDPA reveals about students’ understanding of concepts related to uncertainty during their first comprehensive year of introductory lab courses, we now turn to what the CDPA can tell us about the level of learning that takes place in more advanced laboratory courses. As previously explained, beyond the 200-level courses, this study also includes data collected at the 300- and 400-levels, most notably PHYS 359 - Advanced Physics Laboratory 1, reserved for Honors Physics undergraduate students, and PHYS 439 - Majors Laboratory in Modern Physics, the equivalent course for Physics majors.
@@
Between 2019 and 2022, the variability in CDPA score for PHYS 258 demonstrates that, at this point in their learning, students have not mastered the skills needed to evaluate experimental uncertainties. For PHYS 359, however, we note a significant increase in the CDPA score in 2022 that warrants a more in-depth investigation into what factors may have contributed to this leap in student understanding. As can be seen in Figure 2, the 2022 mean CDPA score was significantly higher, from a class average of $2.5\pm 2.0$ by the end of PHYS 258 to $4.5\pm 1.0$ by the end of PHYS 359. Although we can expect noticeable improvement in the CDPA score after this upper-level laboratory course, the extent of this statistically significant increase over one semester left us with more questions.
@@
Through discussion with the instructor who taught PHYS 359 during the Winter 2022 semester, we could ascertain that certain structural elements of the course contributed to this demonstrable increase in student understanding of uncertainties. Three key factors stand out in the course design and context that semester. First, the most significant difference between PHYS 258 and PHYS 359 is the class size. While PHYS 258 had over 100 students, the smaller enrollment numbers in PHYS 359 allowed for greater one on one interaction between students and their instructor. Second, students also spent more time in the lab each week, and third, they performed experiments that require a higher degree of precision and attention to uncertainties.
@@
The PHYS 359 course has undergone changes and adaptations throughout the years. One example is that a course instructor for PHYS 359 in 2022 identified a noticeable weakness in incoming students’ level of understanding of uncertainties. This led to the development and implementation of an additional module dedicated to curve fitting in a self-paced Python tutorial. This module offered students more advanced exercises emphasizing graph interpretation and analysis, incorporating an interactive simulation feature and enabling personalized feedback. The teaching tool allowed the instructor to leverage the smaller class size for a more individualized instructional approach, dedicating time to helping students apply concepts related to uncertainties. With this personalized tool, the instructor could identify and correct student misconceptions more precisely.
@@ V Conclusions
By comparing average CDPA scores from pre and post tests across all core laboratory courses in the McGill experimental physics undergraduate curriculum, we were able to distinguish distinct patterns in students’ understanding of concepts related to uncertainties from introductory 100-level courses through upper 300-level courses. As a general observation, our results show that, at the 100-level, students’ CDPA scores remained quite low and were mostly on-par with random guesswork. These results confirm UBC’s findings for introductory-level courses, which can be explained by the fact that, in first-year courses, students are introduced to the concept of uncertainties for the first time and are only expected to solve simple problems and conduct experiments without the need of deep conceptual understanding of uncertainties.
@@
Across all four years of this study, we noticed an increase in students’ CDPA score from the 100 to the 200 level. This uptick in student understanding may be explained by the fact that at the 200-level, students complete a full academic year of laboratory courses, where they learn curve fitting, consideration of uncertainties, and production and interpretation of graphs. In PHYS 257 students are asked to assess systematic errors and perform error propagation and in PHYS 258 they need to consider statistical uncertainties to get the best-fit curves, understand and justify their reasoning, and differentiate between different models. In addition, PHYS 258 culminates in a small research project in which students are required to answer a research question of their choice, experimentally validate their measurements and demonstrate understanding of how the uncertainties should be considered in the analysis. Nevertheless, it is not until the 300-level and 400-level courses, when 20 percent of the lab report grade is allocated to justifying quantitative statistical uncertainties, that students’ demonstrated understanding of uncertainties develops into holistic concepts. In addition to using the CDPA instrument as a mean to rank student understanding across the departmental laboratory curricula, our research also shows that this tool can be used to help professors comprehend the nature of their students’ misconceptions. As shown in the breakdown of CDPA incorrect responses, Figure 4, CDPA results can be used as a remediation tool to help professors identify gaps in student understanding, thereby providing an additional means to refine their teaching approach for subsequent cohorts.
@@
In conclusion, through this collaborative study, we were able to highlight new opportunities for pedagogical and structural improvement in physics laboratory courses. We show how a better understanding of students misconceptions can enhance professors’ understanding of student learning on a course-by-course basis. This research reinforces the value of collecting program level data using a common instrument and then reporting the findings back to course professors for interpretation. It is important to note that engaging the professors in the process of data analysis, allowed us to pinpoint the educational benefit featured of this research. This study also points to a potential need for more inquiry-guided learning in laboratory courses, or, at the very least, it serves as an impetus to carve out more personalized attention in large-scale introductory laboratory courses.